x^2+8x=9 is there any way to make this understandable. I am really confused on how to solve these problems!!!
Answer by Alan3354(69443)

x^2+8x=9 is there any way to make this understandable. I am really confused on how to solve these problems!!!
-------------
There's more than one way. The quadratic equation ALWAYS works, if you know about it.
It's quicker and easier if you can factor it, tho.
x^2+8x=9
x^2 + 8x - 9 = 0 (always get it equal to zero)
(x + 9)*(x - 1) = 0
If the product is zero, either (x+9) is zero, or (x-1) is zero.
x = -9
x = 1